OpenStudy (anonymous):

Select the equations that are parallel and perpendicular to y = 2x - 1 and that pass through the point (4, 1). parallel: y = 2x - 7 perpendicular: y = negative one over twox + 3 parallel: y = 2x + 2 perpendicular: y = negative one over twox + four and one half parallel: y = negative one over twox - 7 perpendicular: y = 2x + 3 parallel: y = -2x - 7 perpendicular: y = negative one over two x - 1

OpenStudy (anonymous):

If y=mx+c line parallel to this line is y=mx+c1,find c1 from this eq. by the given point and substitute in this eq. slope of line perpendicular to y=mx+c is -1/m Hence line perpendicular to y=mx+c is y=(-1/m)x+c2,find c2 by the given condition and substitute in this eq.

OpenStudy (anonymous):

I'm so confused

OpenStudy (issy14):

Parallel lines have equal slope. Perpendicular lines have a negative reciprocal of = -1 Select the equations that are parallel and perpendicular to y = 2x - 1 and that pass through the point (4, 1). Point:(4,1) y= 2x-1 m(slope) = 2 1- point slope formula= y-y1=m(x-x1) y-1= 2 (x-4) y-1=2x-8 y=2x-8+1 y=2x-7 Parallel: y = 2x - 7 This is parallel = the slope is 2, the same as the equation. but it has the same y- intercept so basically is the same line. Perpendicular: y = -1/2x + 3 m= -1/2 -1/2 : 2= -2/2 = -1 This one is perpendicular and it has a different y-intercept. Parallel: y = 2x + 2 This one has the same slope =2. And different y-intercept. Perpendicular: y = -1/2x + 4.5 This one is perpendicular. This is your answer. Parallel: y = negative one over twox - 7 Perpendicular: y = 2x + 3 Not perpendicular, the reciprocal of 2 : 2 = 1 not -1. Parallel: y = -2x - 7 Not parallel, it doesn't have the same slope. Perpendicular: y = -1/2x - 1